OpenStudy (anonymous):

simplify by removing factors of 1 s^2 - 49 / s^2 - 14s + 49 = the simplified form is __

OpenStudy (anonymous):

simplify by removing factors of 1 s^2 - 49 / s^2 - 14s + 49 = the simplified form is __

OpenStudy (anonymous):

(s+7)/(s-7)

OpenStudy (anonymous):

s^2 - 49 / s^2 - 14s + 49 = ((s-7)(s+7))/((s-7)(s-7))=((s-7)/(s-7))*(s+7)/(s-7) s^2 - 49 / s^2 - 14s + 49 = (s+7)/(s-7)